Qno:01


In Pakistan internet service is ought to be considered as the perfectly competitive industry. Following is the information available for the internet service provider A operating in the market. 

 

Total output

Total Cost


180

10

280

20

400

30

530

40

670

50

820

60

980

70

1150

 

a. Briefly discuss the characteristics of the market Service Provider A is operating.​2 Marks

b. If the market price of the fruit is 16​​​​​​6 Marks

 

i. Find how much output will the firm produce (Show calculations).​2 Marks

ii. Illustrate it graphically​​​​​​​2 Marks

iii. Find if the firm is maximizing profit or minimizing loss (show calculations) and highlight the same in the graph drawn in part ii. ​​​​​1 Mark

iv. From your answer in part iii, explain if the firm is operating in the LR or SR. 1 Mark


1
Expert's answer
2021-06-01T14:00:09-0400

The characteristics of a perfectly competitive market where provider A operates are listed below;

  • There are many buyers and sellers in the perfect competition.
  • Every firm sells homogenous products.
  • There are no barriers to entry and exit.
  • The firm is a price taker. In the perfectly competitive industry firms are price takers and sell whatever quantity they want at the market price fixed by the industry.
  • Marginal revenue in this market is equal to the price. It is because the price is fixed and when an additional unit is sold the revenue received is equal to the price itself.
  • Buyers and sellers are informed about the products.


(b).

Given,

Market price = 16

(1).

"Marginal \space cost =\frac{change\space in\space total\space cost}{change\space in\space output}"





The marginal revenue is equal to the marginal cost at the output level of 60. So the firm will produce 60 units.


the x-axis measures the quantity of output, and the y-axis measures the price and marginal revenue. The MC curve intersects the MR curve at point E where the quantity is 60 units. So the firm will produce 60 units. At equilibrium, the MC curve should be rising (upward sloping). At point E the MC curve intersects while rising so it is the equilibrium point.



(2).

"Average\space cost=\frac{Total\space cost}{Total\space output}"


To find the profit or loss we also need average cost and total revenue.





According to the table, at the equilibrium level of output, the average cost is greater than the price (16.33>16). So the firm incurs a loss.

Loss = Total cost - Total revenue

Loss = 980 - 960

Loss = 20

So the firm is minimizing loss.

Graphical presentation:




According to the above figure, the area between the price line and the AC at the output level of 60 shows the loss. The yellow area shows the loss.


(3). 

The firm is operating in the short run. It is because, in the long run, a firm earns a normal profit. At a normal profit, the price is equal to the price. But here the price is lower than the AC. So the firm is operating in the short run.